Assume the credit terms offered to your firm by your suppliers are 2.9​/5​, Net 30. Calculate the cost of the trade credit if your firm does not take the discount and pays on day 30.

The effective annual cost of the trade credit is _____%. ​(Round to two decimal​ places.)
Answer #1

Solution:

The formula for calculating the effective annual cost of trade credit if the firm does not take the discount is

= ( 1 + [Discount rate / (1 – Discount Rate )] ) n – 1

Where n= No. of days in a period / ( Total payment Period – period for which discount has been

offered)

As per the information given in the question we have

Discount rate = 2.9 % = 0.029   ; No. of days in a period = 365 days ;

Total payment Period = 30 days ; Period for which discount has been offered = 5 days

Total payment Period – period for which discount has been offered = 30 – 5 = 25 days

Applying the above values in the formula we have

= ( 1 + (0.029 / (1 – 0.029)) ( 365 / 25 )   - 1                        

= ( 1 + (0.029 / 0.971 )) (14.6)   - 1                                    

= ( 1 + 0.029866 ) ( 14.6 )   - 1                                             

= ( 1.029866 ) ( 14.6 )   - 1

= 1.5367 – 1 = 0.5367   = 53.67 %                                                

Thus the annual effective cost of credit is = 53.67 % ( when rounded off to two decimal places )  

Note: ( 1.029866 ) ( 14.6 )   is calculated using the excel formula =POWER(Number,Power)

=POWER(1.029866,14.6) = 1.5367